OpenStudy (iluvsoccer):

Which circles have an area of 16pi in^2? Choose all answers that are correct. A. a circle with a radius of 2 in. B. a circle with a radius of 4 in. C. a circle with a diameter of 8 in. D. a circle with a diameter of 16 in.

OpenStudy (sepeario):

Do you know the formula for the area of a circle?

OpenStudy (iluvsoccer):

Yes Mr. Bean. xD :P No offense. Just pulling your legs.

OpenStudy (sepeario):

Ok what gives us 16 when we square it?

OpenStudy (iluvsoccer):

8

OpenStudy (sepeario):

Think again.. 8 squared doesn't equal 16.

OpenStudy (iluvsoccer):

4.

OpenStudy (anonymous):

o3o yep your right 4x4=16 ^~^

OpenStudy (anonymous):

Sorry. I switched accounts. That account is lagging.

OpenStudy (anonymous):

area = pi r^2 pi r^2=16 pi r^2=16 r= sqrt(16)=4

OpenStudy (iluvsoccer):

@jabez177

jabez177 (jabez177):

Awesome job!

jabez177 (jabez177):

Thanks for your help!